# Quantum correlations and metrological advantage among Unruh-DeWitt detectors in de Sitter spacetime

### v1

@@ Abstract
A long-standing debate on Gibbons-Hawking (GH) decoherence centers on its unclear thermal nature. In this work, we investigate the robustness of quantum Fisher information (QFI) and local quantum uncertainty (LQU) in the presence of GH decoherence, using free-falling Unruh-DeWitt (UDW) detectors in de Sitter spacetime (dS-ST). The UDW detectors interact with a massless scalar field in dS-ST and are modeled as open quantum systems, with the field acting as the environment for which we use a master equation to describe their evolution. Our analysis investigates the roles of energy spacing, GH temperature, initial state preparation, and various de Sitter-invariant vacuum sectors on the optimization of QFI and LQU. We find that the optimal values of QFI and LQU depend on the selected de Sitter-invariant vacuum sector and increase with larger energy spacing. Our findings reveal that QFI exhibits resilience to GH decoherence, maintaining a pronounced local peak across a wider range of parameters. This robustness can be further enhanced through strategic initial state preparation and increased energy spacing, resulting in a higher maximum QFI value even under significant environmental decoherence. Our results underscore the critical role of GH thermality in governing QFI and LQU, offering valuable insights for advances in relativistic quantum metrology (RQM).
@@ I Introduction
The Unruh-DeWitt (UDW) detector is a crucial tool for probing quantum fields within specific spacetime geometries [21, 10]. This detector is a microscopic two-level system (qubit) that couples locally to fluctuating backgrounds. In dS-ST, a co-moving UDW detector detects radiation with a thermal spectrum defined by the temperature $T_{GH}=\frac{H}{2\pi}$, where $H$ is the Hubble parameter [11]. This phenomenon, known as the GH effect, demonstrates universality in various quantum fields. For a massless scalar field [22], the infrared divergence of the propagator has negligible influence, leaving the detector’s overall response largely unaffected. To explore the dynamic properties of quantum fields in CSTs [10], the open quantum systems framework has gained prominence [23, 24]. This approach has proven valuable in examining the behavior of UDW detectors in diverse spacetime backgrounds [25, 26, 27]. Within this framework, the UDW detector is modeled as a localized open quantum system, while the fluctuations of the background quantum field act as an environment that induces dissipation and dephasing.
@@
In quantum estimation theory, a central challenge is estimating unobservable parameters in a labeled quantum system using measurement data [43, 44]. Enhancing the precision of these estimates and achieving the quantum accuracy limit are key objectives. Due to its practical relevance, the quantum Fisher information (QFI) has been explored from multiple perspectives, leading to significant progress. For instance, it has been used to establish a statistical generalization of the Heisenberg uncertainty principle [45]. The measurement of QFI with finite precision was demonstrated in [46], and its connections to quantum correlations (QCs) [47, 48] and quantum coherence [49] have been established.
Recently, there has been a growing interest in integrating quantum metrology and quantum information within relativistic frameworks, particularly in exploring quantum fields in CSTs [50, 4, 51]. Additionally, numerous studies have investigated the characteristics of QFI in non-inertial frames [52, 53]. By leveraging relativistic effects in quantum systems, it has shown that measurement precision can be greatly improved [54]. This emerging discipline, referred to as relativistic quantum metrology (RQM), seeks to identify the fundamental limits of measurement accuracy when quantum and relativistic effects are jointly considered. Significant relativistic phenomena explored in this field include the Unruh-Hawking effect [55, 56], the expansion rate of Robertson-Walker universes [57], and the properties of Schwarzschild spacetime [58].
@@ I.1 Contribution
Working within the framework of detector-field interactions, we study a localized system of two UDW detectors in dS-ST. Assuming weak interaction with a bath of scalar fields, the system behaves like an open quantum system, experiencing environmental decoherence due to the thermal nature of dS-invariant vacua, such as the GH effect. In this paper, we evaluate the LQU and QFI at the final equilibrium state of the system and analyze how factors such as the detector’s energy spacing, GH decoherence, the choice of initial state, and the selection of $\alpha$-vacua influence LQU and QFI. Our findings highlight the robustness and enhanced sensitivity of LQU and QFI, illustrating their potential to enhance quantum estimation techniques in practical applications.
@@ I.2 Organization
The rest of the paper is structured as follows: Sec. II introduces the model, where two UDW detectors interact with a scalar field in a dS-ST background. The master equation in Lindblad form is solved to obtain the final equilibrium state. This section also examines fundamental concepts, including QFI and LQU for bipartite states. In Sec. III, we explore the dynamics of UDW in non-BD vacua, which depend on the initial state preparation of the detectors, the GH decoherence effect, and different choices of superselection sectors of $\alpha$-vacua. Sec. IV presents our results, and Sec. V concludes the paper.
@@ II Preliminaries
This section presents the essential concepts for defining QFI and LQU for a localized UDW in dS-ST. We begin by describing the dynamics of a bipartite UDW, governed by a Lindblad master equation. Next, we examine the definitions and properties of QFI and LQU for bipartite systems.
@@ II.1 Markovian dynamics of UDW detectors
To proceed, we should first explore the Markovian evolution of an accelerating UDW detector interacting with a scalar field. This interaction can be modeled as an open quantum system weakly coupled to a heat bath. The total Hamiltonian of the system, which includes the UDW detectors and the environment, is given by
@@
Here, $\mathcal{H}_{D}$ represents the Hamiltonian of two independent detectors in the common co-moving frame. The internal dynamics of each two-level detector is governed by a matrix $2\times 2$, and $\mathcal{H}_{D}$ can be concisely expressed as [5, 23, 27, 4].
@@
The master equation (4) enables us to describe the asymptotic equilibrium states of detectors at large times, which are governed by the competition between environment dissipation in CST background and QC generated through the Markovian dynamics of detectors [25, 62]. For two-detector system, the initial interatomic separation $L\equiv\left|\textbf{x}^{(a)},\textbf{x}^{(b)}\right|$ is a control parameter of correlation generation, as Kossakowski matrices now become distance-dependent since in general $\mathcal{Y}^{(ab)}=\mathcal{Y}^{(ba)}\equiv\mathcal{Y}(\omega,L)=\mathcal{Y}_{0}(\omega)f(\omega,L)$ for two separated detectors [23, 63], where $f(\omega,L)$ is an even function of frequency $\omega$. It is not surprising [64] that the correlation generation between detectors would be more effective for smaller $L$, and becomes impossible for an infinitely large separation. It was shown in [65] that a proper $L$ always exists below which the generated correlation can persist asymptotically in final equilibrium states under environment dissipation. Therefore, we can concisely fix a small interatomic separation, and only be concerned about the influence of environment decoherence on the equilibrium states of detectors. In such a situation, all the Kossakowski matrices become equal $\Omega_{ij}^{(aa)}=\Omega_{ij}^{(bb)}=\Omega_{ij}^{(ab)}=\Omega_{ij}^{(ba)}$ [26], where
@@
After solving the master equation (4), the final reduced density matrix of two-detector at asymptotic equilibrium can be expressed in a Bloch form [25, 62]
@@
Here, the ratio $T=\frac{\kappa_{-}}{\kappa_{+}}$ is determined by the dynamics of the system. The final equilibrium state is also dependent on the choice of initial state by $\tau=\sum_{ii}\eta_{ii}(0)$ which is a constant of motion and satisfies $\tau\in[-3,1]$ to keep $\eta_{ab}(0)$ positive.
@@ II.2 QFI
Here, we will review the concept of QFI and its calculation based on the spectral decomposition of the density operator. We define QFI using a probe state represented by $\eta_{\nu}$, where $\nu$ is an unobservable parameter. The QFI concerning $\nu$ can be formulated as [66]
@@
It is evident that $\mathfrak{F}_{\nu}$ is solely dependent on the support set of $\eta_{\nu}$ and remains unaffected by eigenstates outside this support set. The first term in Eq. (17) depends only on the eigenvalues of $\eta_{\nu}$ and corresponds to the classical part, while the second and third term corresponds to the quantum part. Given the eigenstates outlined in (32), all the vectors $\left|\psi_{i}^{\prime}\right\rangle=0$. Therefore, the QFI is entirely driven by the classical contribution
@@ II.3 LQU
As a key measure in our analysis, we introduce LQU as the minimum skew information obtainable from a single local measurement. Let $\eta_{ab}$ be the state of a bipartite system. Define $K_{\Delta}=K_{\Delta}^{(a)}\otimes\hat{s}_{0}^{(b)}$ as a local observable, where $K_{\Delta}^{(a)}$ is a Hermitian operator on subsystem $(a)$ with spectrum $\Delta$, and $\hat{s}_{0}^{(b)}$ is the identity operator acting on subsystem $(b)$. We assume that $\Delta$ is non-degenerate, corresponding to maximally informative observables on subsystem $(a)$. The LQU with respect to subsystem $(a)$, optimized over all local observables on $(a)$ with non-degenerate spectrum $\Delta$, is expressed as [68, 69].
@@
Based on this Ref-[70], we summarize the key characteristics of skew information: it is nonnegative, equals zero if and only if the state and the observable commute, and is convex, meaning it decreases under classical mixing. If a local operator $K_{\Delta}$ exists such that $\Upsilon(\eta_{ab},K_{\Delta})=0$, then the quantum system described by the state $\eta_{ab}$ shows no QC between its two subsystems. For bipartite states the LQU is expressed as [69]
@@ III QFI and LQU in non-BD vacua
The dS-ST represents the simplest solution of Einstein’s field equations when a positive cosmological constant, $\Lambda$, is present. dS-ST is maximally symmetric, characterized by the isometry group $\mathrm{SO}(4,1)$ in four dimensions. Its importance arises from two primary aspects. Firstly, the observed accelerated expansion of our Universe strongly suggests that it is influenced by a small, positive $\Lambda$ or an equivalent form of dark energy. Secondly, the remarkable uniformity and isotropy of the Universe on large scales, as observed in the CMB, imply that the early Universe likely underwent a phase of rapid accelerated expansion, commonly referred to as cosmic inflation for which we suggest the reader see the Ref-[71] for further details.
@@
$Y^{+}_{BD}(x(t),y(t^{\prime}))=-\frac{H^{2}}{16\pi^{2}\left[(t-t^{\prime})H/2-i\epsilon\right]}$ (23)
@@
By Mottola–Allen transformation, one can further define a one-parameter family of de Sitter-invariant vacua called $\alpha-$vacua $\left|\alpha\right\rangle$, each of which can be interpreted as a squeezed state over $\left|BD\right\rangle$, i.e., $\left|\alpha\right\rangle=\hat{S}(\alpha)\left|BD\right\rangle$, where $Re\alpha<0$ and $\hat{S}(\alpha)$ denotes a squeezing operator in quantum optics [37, 35, 36, 34, 33]. In particular, we adopt CPT invariant $\alpha-$vacua, which means $\alpha<0$ is real. Therefore, the Wightman function for the scalar field in $\alpha-$vacua can be expressed in terms of $Y^{+}_{BD}$ as
@@
For fixed energy spacing of detectors, it is easy to find that $T_{\alpha}\in\left[-1,+1\right]$ as $\beta$ varying. Inserting (27) into (11), we obtain the final equilibrium state of two detectors respecting to $\alpha-$vacua, which is
@@
We choose to estimate the GH parameter $\beta$ for any arbitrary initial state. The related $\mathfrak{F}_{\beta}$ can be straightforwardly calculated from (17), by substituting the eigenvalues (31) of the diagonalized density matrix of detectors system (28). According to different preparations of initial state encoded in $\tau\in[-3,1]$, we come to two classes of QFI:
(i) $\tau=-3$, only one nonvanishing eigenvalue $\mu_{4}=1$, which gives $\mathfrak{F}_{\beta}=0$;
@@
The explicit form of the matrix elements, which depend on the system parameters from (28) and the GH decoherence environment, is derived as
@@
For $\tau\in[-3,1]$, the situation for (33) and (3435) is complicated due to its dependence on the infinite family of non-BDV. In the following, our aim is to demonstrate that for certain initial states, QCs can be generated in the final state (28) after the Markovian evolution of the system. To illustrate this, we consider two UDW detectors initially prepared in a Bell-diagonal state on a freely falling basis
@@
with coefficients $0\leq|r_{i}|\leq 1$. This state is the convex combination of four Bell states and reduce to maximally entangled states (Bell-basis) if $|r_{1}|=|r_{2}|=|r_{3}|=1$.

### v2

@@ Abstract
A long-standing debate on Gibbons-Hawking (GH) decoherence centers on its obscure thermal nature. In this work, we investigate the robustness of quantum Fisher information (QFI) and local quantum uncertainty (LQU) in the presence of GH decoherence, using free-falling Unruh-DeWitt (UDW) detectors in de Sitter spacetime (dS-ST). The UDW detectors interact with a massless scalar field in dS-ST and are modeled as open quantum systems, with the field serving as the environment, described by a master equation that outlines their evolution. Our analysis investigates the roles of energy spacing, GH temperature, initial state preparation, and various de Sitter-invariant vacuum sectors on the optimization of QFI and LQU. We find that the optimal values of QFI and LQU depend on the selected de Sitter-invariant vacuum sector and increase with larger energy spacing. Our findings reveal that QFI exhibits resilience to GH decoherence, maintaining a pronounced local peak across a broader range of parameters. This robustness can be further enhanced through strategic initial state preparation and increased energy spacing, resulting in a higher maximum QFI value even under significant environmental decoherence. Our results underscore the critical role of GH thermality in governing QFI and LQU, offering valuable insights for advances in relativistic quantum metrology (RQM).
@@ I Introduction
The Unruh-DeWitt (UDW) detector is a crucial tool for probing quantum fields within specific spacetime geometries [21, 10, 22, 23]. This detector is a microscopic two-level system (qubit) that couples locally to fluctuating backgrounds. In dS-ST, a co-moving UDW detector detects radiation with a thermal spectrum defined by the temperature $T_{GH}=\frac{H}{2\pi}$, where $H$ is the Hubble parameter [11]. This phenomenon, known as the GH effect, demonstrates universality in various quantum fields. For a massless scalar field [24], the infrared divergence of the propagator has negligible influence, leaving the detector’s overall response largely unaffected. To explore the dynamic properties of quantum fields in CSTs [10], the open quantum systems framework has gained prominence [25, 26]. This approach has proven valuable in examining the behavior of UDW detectors in diverse spacetime backgrounds [27, 28, 29, 30]. Within this framework, the UDW detector is modeled as a localized open quantum system, while the fluctuations of the background quantum field act as an environment that induces dissipation and dephasing.
@@
In quantum estimation theory, a central challenge is estimating unobservable parameters in a labeled quantum system using measurement data [46, 47]. Enhancing the precision of these estimates and achieving the quantum accuracy limit are key objectives. Due to its practical relevance, the QFI has been explored from multiple perspectives, leading to significant progress. For instance, it has been used to establish a statistical generalization of the Heisenberg uncertainty principle [48]. The measurement of QFI with finite precision was demonstrated in [49], and its connections to quantum correlations (QCs) [50, 51] and quantum coherence [52] have been established. Recently, there has been a growing interest in integrating quantum metrology and quantum information within relativistic frameworks, particularly in exploring quantum fields in CSTs [53, 4, 54]. Furthermore, numerous studies have investigated the characteristics of QFI in non-inertial frames [55, 56]. By leveraging relativistic effects in quantum systems, it has shown that measurement precision can be greatly improved [57]. This emerging discipline, referred to as RQM, seeks to identify the fundamental limits of measurement accuracy when quantum and relativistic effects are jointly considered. Significant relativistic phenomena explored in this field include the Unruh-Hawking effect [58, 59], the expansion rate of Robertson-Walker universes [60], and the properties of Schwarzschild spacetime [61].
@@ I.1 Contribution
Working within the framework of detector-field interactions, we study a system of two UDW detectors in dS-ST. Assuming a weak interaction with a bath of scalar fields, the system behaves like an open quantum system, experiencing environmental decoherence due to the thermal nature of dS-invariant vacua, such as the GH effect. In this paper, we evaluate the LQU and QFI at the final equilibrium state of the system and analyze how factors such as the detector’s energy spacing, GH decoherence, the choice of initial state, and the selection of $\alpha$-vacua influence LQU and QFI. Our findings highlight the robustness and enhanced sensitivity of LQU and QFI, illustrating their potential to enhance quantum estimation techniques in practical applications.
@@ I.2 Organization
The remainder of the paper is structured as follows. Sec. II introduces the model, in which two UDW detectors interact with a scalar field in a dS-ST background. The master equation in Lindblad form is solved to obtain the final equilibrium state. This section also examines fundamental concepts, including QFI and LQU for bipartite states. In Sec. III, we explore the dynamics of UDW in non-BD vacua, which depend on the initial state preparation of the detectors, the GH decoherence effect, and different choices of superselection sectors of $\alpha$-vacua. Sec. IV presents our results, and Sect. V concludes the paper.
@@ II Preliminaries
This section presents the essential concepts for defining QFI and LQU for a localized UDW in dS-ST. We begin by describing the dynamics of a bipartite UDW governed by a Lindblad master equation. Next, we examine the definitions and properties of QFI and LQU for bipartite systems.
@@ II.1 Markovian dynamics of UDW detectors
To proceed, we should first explore the Markovian evolution of an accelerating UDW detector interacting with a scalar field. The total Hamiltonian of the system, which includes the UDW detectors and the environment, is given by
@@
Here, $\mathcal{H}_{D}$ represents the Hamiltonian of two independent detectors in the standard co-moving frame. The internal dynamics of each two-level detector is governed by a matrix $2\times 2$, and $\mathcal{H}_{D}$ can be concisely expressed as [5, 25, 29, 4].
@@
The master equation (4) enables us to describe the asymptotic equilibrium states of detectors at large times, which are governed by the competition between environment dissipation in CST background and QC generated through the Markovian dynamics of detectors [27, 65]. For a two-detector system, the initial interatomic separation $L\equiv\left|\textbf{x}^{(a)},\textbf{x}^{(b)}\right|$ is a control parameter of correlation generation, as Kossakowski matrices now become distance-dependent since, in general, $\mathcal{Y}^{(ab)}=\mathcal{Y}^{(ba)}\equiv\mathcal{Y}(\omega,L)=\mathcal{Y}_{0}(\omega)f(\omega,L)$ for two separated detectors [25, 66], where $f(\omega,L)$ is an even function of frequency $\omega$. It is not surprising [67] that the correlation generation between detectors would be more effective for smaller $L$, and becomes impossible for an infinitely large separation. It was shown in [68] that a proper $L$ always exists below which the generated correlation can persist asymptotically in final equilibrium states under environment dissipation. Therefore, we can concisely fix a small interatomic separation and only be concerned about the influence of environment decoherence on the equilibrium states of detectors. In such a situation, all the Kossakowski matrices become equal $\Omega_{ij}^{(aa)}=\Omega_{ij}^{(bb)}=\Omega_{ij}^{(ab)}=\Omega_{ij}^{(ba)}$ [28], where
@@
After solving the master equation (4), the final reduced density matrix of two detectors at asymptotic equilibrium can be expressed in the Bloch form [27, 65]
@@
Here, the ratio $T=\frac{\kappa_{-}}{\kappa_{+}}$ is determined by the dynamics of the system. The final equilibrium state also depends on the choice of initial state $\tau=\sum_{ii}\eta_{ii}(0)$, which is a constant of motion and satisfies $\tau\in[-3,1]$ to keep $\eta_{ab}(0)$ positive.
@@ II.2 QFI
Here, we review the concept of QFI and its calculation based on the spectral decomposition of the density operator. We define QFI using a probe state represented by $\eta_{\nu}$, where $\nu$ is an unobservable parameter. The QFI concerning $\nu$ can be formulated as [69]
@@
It is evident that $\mathfrak{F}_{\nu}$ is solely dependent on the support set of $\eta_{\nu}$ and remains unaffected by eigenstates outside this support set. The first term in Eq. (17) depends only on the eigenvalues of $\eta_{\nu}$ and corresponds to the classical part, while the second and third terms correspond to the quantum part. Given the eigenstates described in (32), all the vectors $\left|\psi_{i}^{\prime}\right\rangle=0$. Therefore, the QFI is entirely driven by the classical contribution
@@ II.3 LQU
As a key measure in our analysis, we introduce LQU as the minimum skew information obtainable from a single local measurement. Let $\eta_{ab}$ be the state of a bipartite system. Define $K_{\Delta}=K_{\Delta}^{(a)}\otimes\hat{s}_{0}^{(b)}$ as a local observable, where $K_{\Delta}^{(a)}$ is a Hermitian operator on the subsystem $(a)$ with spectrum $\Delta$, and $\hat{s}_{0}^{(b)}$ is the identity operator acting on subsystem $(b)$. We assume that $\Delta$ is non-degenerate, corresponding to maximally informative observables on subsystem $(a)$. The LQU with respect to subsystem $(a)$, optimized over all local observables on $(a)$ with non-degenerate spectrum $\Delta$, is expressed as [71, 72].
@@
Based on this Ref-[73], we summarize the key characteristics of skew information: it is nonnegative, equals zero if and only if the state and the observable commute, and is convex, meaning it decreases under classical mixing. If a local operator $K_{\Delta}$ exists such that $\Upsilon(\eta_{ab},K_{\Delta})=0$, then the quantum system described by the state $\eta_{ab}$ shows no QC between its two subsystems. For bipartite states, the LQU is expressed as [72]
@@ III QFI and LQU in non-BD vacua
The dS-ST represents the simplest solution of Einstein’s field equations when a positive cosmological constant, $\Lambda$, is present. dS-ST is maximally symmetric, characterized by the isometry group $\mathrm{SO}(4,1)$ in four dimensions. Its importance arises from two primary aspects. Firstly, the observed accelerated expansion of our Universe strongly suggests that it is influenced by a small, positive $\Lambda$ or an equivalent form of dark energy. Secondly, the remarkable uniformity and isotropy of the Universe on large scales, as observed in the CMB, imply that the early Universe likely underwent a phase of rapid accelerated expansion, commonly referred to as cosmic inflation, for which we suggest the reader to see the Ref-[74] for further details.
@@
$Y^{+}_{BD}(x(t),y(t^{\prime}))=-\frac{H^{2}}{16\pi^{2}\left[(t-t^{\prime})H/2-i\epsilon\right]},$ (23)
@@
By Mottola–Allen transformation, one can further define a one-parameter family of de Sitter-invariant vacua called $\alpha-$vacua $\left|\alpha\right\rangle$, each of which can be interpreted as a squeezed state over $\left|BD\right\rangle$, i.e., $\left|\alpha\right\rangle=\hat{S}(\alpha)\left|BD\right\rangle$, where $\text{Re}(\alpha)<0$ and $\hat{S}(\alpha)$ denotes a squeezing operator in quantum optics [40, 38, 39, 37, 36]. Throughout this work, we adopt CPT invariant $\alpha-$vacua, which means $\alpha<0$ is real, with $|\alpha|$ denoting the magnitude of $\alpha$ in all figures and analyses. Therefore, the Wightman function for the scalar field in $\alpha-$vacua can be expressed in terms of $Y^{+}_{BD}$ as
@@
For fixed energy spacing of detectors, it is easy to find that $T_{\alpha}\in\left[-1,+1\right]$ as $\beta$ varies. Inserting (27) into (11), we obtain the final equilibrium state of two detectors respecting to $\alpha-$vacua, which is
@@
We choose to estimate the GH parameter $\beta$ for any arbitrary initial state. The related $\mathfrak{F}_{\beta}$ can be straightforwardly calculated from (17), by substituting the eigenvalues (31) of the diagonalized density matrix of the detectors system (28). According to different preparations of the initial state encoded in $\tau\in[-3,1]$, we come to two classes of QFI:
(i) $\tau=-3$, only one non-vanishing eigenvalue $\mu_{4}=1$, which gives $\mathfrak{F}_{\beta}=0$;
@@
The explicit form of the matrix elements (21), which depend on the system parameters from (28) and the GH decoherence environment, is derived as
@@
For $\tau\in[-3,1]$, the situation for (33) and (3435) is complicated due to its dependence on the infinite family of non-BDV. In the following, our aim to demonstrate that for certain initial states, QCs can be generated in the final state (28) after the Markovian evolution of the system. To illustrate this, we consider two UDW detectors initially prepared in a Bell-diagonal state on a freely falling basis
@@
with coefficients $0\leq|r_{i}|\leq 1$. This state is the convex combination of four Bell states and reduces to maximally entangled states (Bell-basis) if $|r_{1}|=|r_{2}|=|r_{3}|=1$.